OpenStudy (anonymous):

Verify the identity. (1 + tan^2u)(1-sin^2u) =1

OpenStudy (anonymous):

We have: \[ (1+\tan^2(u))(1-\sin^2(u))=1 \]We know: \[ 1+\tan^2(u)=\sec^2(u) \]And: \[ \sin^2(u)+\cos^2(u)=1\iff\\ 1-\sin^2(u)=\cos^2(u) \]So: \[ (\sec^2(u))(\cos^2(u))=\frac{\cos^2(u)}{\cos^2(u)}=1 \]